# Semantic Reasoning Meets Numerical Precision: An LLM-Powered Multi-Agent System for Power Grid Control

### v1

@@ V-A Test Network Configuration
We validated Grid-Agent’s performance on a diverse set of standard IEEE and CIGRE test systems to ensure the framework’s applicability across networks of varying size, topology, and complexity. The specific test cases were generated by introducing multiple, often interacting, violations to stress-test the agent’s reasoning and coordinated control capabilities. The networks used include:
  • IEEE 30-Bus System: We created two scenarios based on this transmission system model:
    • Case30 Light: A baseline scenario with 2 voltage and 3 thermal violations.
    • Case30 Medium: A simpler thermal violation scenario used to test focused problem-solving.
  • CIGRE Medium Voltage (MV) Distribution Network: this 15-bus European benchmark network was used to simulate severe distribution-level challenges:
    • CIGRE MV Severe: A highly stressed scenario with 14 concurrent voltage and thermal violations.
    • CIGRE MV Disconnected: A contingency scenario where a line fault isolates 5 buses from the main grid, testing the agent’s ability to restore connectivity.
  • IEEE 69-Bus Distribution System: This larger radial distribution network was used to evaluate scalability and performance in more complex topologies:
    • IEEE 69 Large Loads: A scenario with 18 simultaneous voltage and thermal violations due to heavy loading.
    • IEEE 69 Medium Loads: A highly complex case with 29 violations, representing a significant system disturbance.
    • IEEE 69 Disconnected: A fault scenario resulting in 13 disconnected buses, requiring topological reconfiguration.
@@ V-B LLM Test Suite Configuration
The core of our study involved comparing the performance of the Grid-Agent framework when powered by different LLMs. We selected six representative models to cover a spectrum of capabilities and sizes: gemini-2.5-pro, gemini-2.5-flash, gemini-2.5-flash-lite, gpt-4.1, gpt-4.1-mini, and gpt-4.1-nano. Each LLM was integrated into the Planner Agent role within the same multi-agent architecture. The performance of these configurations was compared with each other to identify the most effective reasoning engines for this task. Table II summarizes the configurations of each LLM model used in the experiment.
@@
gpt-4.1 OpenAI Temperature: 0.1
gpt-4.1-mini OpenAI Temperature: 0.1
gpt-4.1-nano OpenAI Temperature: 0.1
@@
gemini-2.5-flash Google Standard Config
gemini-2.5-flash-lite Google Standard Config
@@ V-C Performance Metrics
System performance was evaluated using a suite of quantitative metrics designed to capture the effectiveness, efficiency, and quality of the violation resolution process:
  • Success Rate: The primary metric, defined as the percentage of violation scenarios successfully and fully resolved.
  • Action Efficiency: The average number of violations resolved per control action, measuring the agent’s ability to formulate coordinated and impactful solutions.
  • Convergence Speed: The average number of planning-execution-validation iterations required for complete resolution.
  • Solution Quality: A qualitative measure of the final solution, prioritizing the minimization of total actions and avoiding disruptive controls like load curtailment where possible.
  • Runtime: The average wall-clock time in seconds to reach a final solution.
@@ V-D Experimental Results and Analysis
As shown in Fig.2a, the more capable and efficient models, gemini-2.5-flash and gpt-4.1-mini, achieved a perfect 100% success rate across all scenarios. The flagship models, gemini-2.5-pro and gpt-4.1, also performed robustly with an 81% success rate. The smaller models, gemini-2.5-flash-lite and gpt-4.1-nano, struggled with the more complex scenarios, achieving 63% and 44% success rates, respectively.
@@ VII Conclusion
Key achievements of this work include the development of a modular multi-agent architecture for decomposed problem-solving, an adaptive multi-scale representation framework that enables LLM application to larger networks, and a coordinated action optimization strategy that resolves multiple violations with minimal intervention. The integrated safety mechanisms, including sand-boxed execution and automated rollbacks, ensure that all proposed solutions are validated before affecting the system, making the framework suitable for critical infrastructure applications.
@@ Acknowledgment
The authors acknowledge the support of the open-source community, particularly the developers of pandapower and LangGraph frameworks that enabled this research.

### v2

@@ V-A Test Network Configuration
We validated Grid-Agent’s performance on a diverse set of standard IEEE and CIGRE test systems to ensure applicability across networks of varying size, topology, and complexity. The specific test cases were generated by introducing multiple, often interacting, violations to stress-test the agent’s reasoning and coordinated control capabilities.
For the IEEE 30-Bus System, two transmission scenarios were created: Case30 Light, a baseline with 2 voltage and 3 thermal violations, and Case30 Medium, a simpler thermal violation scenario.
To simulate severe distribution-level challenges, we used the CIGRE Medium Voltage (MV) Distribution Network, a 15-bus European benchmark, featuring the CIGRE MV Severe scenario with 14 concurrent violations and the CIGRE MV Disconnected scenario, where a fault isolates 5 buses.
Finally, to evaluate scalability, the larger IEEE 69-Bus Distribution System was tested with scenarios including IEEE 69 Large Loads (18 simultaneous violations), IEEE 69 Medium Loads (a highly complex case with 29 violations), and IEEE 69 Disconnected (a fault resulting in 13 disconnected buses).
@@ V-B LLM Test Suite Configuration
The core of our study involved comparing the performance of the Grid-Agent framework when powered by different LLMs. We selected six representative models to cover a spectrum of capabilities and sizes: gemini-2.5-pro, gemini-2.5-flash, gemini-2.5-flash-lite, gpt-4.1, gpt-4.1-mini, and gpt-4.1-nano. Each LLM was integrated into the Planner Agent role within the same multi-agent architecture. The performance of these configurations was compared with each other to identify the most effective reasoning engines for this task. To prioritize inference speed, a critical factor for real-time applications, the thinking budget for the gemini models was minimized. The specific configurations for each model are detailed in Table II.
@@
gpt-4.1 OpenAI Default Config
gpt-4.1-mini OpenAI Default Config
gpt-4.1-nano OpenAI Default Config
@@
gemini-2.5-flash Google Thinking Budget: 0
gemini-2.5-flash-lite Google Thinking Budget: 0
@@ V-C Performance Metrics
System performance was evaluated using a suite of quantitative metrics designed to capture the effectiveness, efficiency, and quality of the violation resolution process. The success rate served as the primary metric, defined as the percentage of violation scenarios that were successfully and fully resolved.
To assess the efficiency of control actions, we measured action efficiency, defined as the average number of violations resolved per control action. This metric reflects the agent’s ability to formulate coordinated and impactful solutions rather than relying on redundant or scattered decisions. Another important aspect was convergence speed, which captures the average number of planning–execution–validation iterations required to achieve complete resolution.
Beyond efficiency and speed, we also considered solution quality. This metric emphasizes not only the resolution of violations but also the manner in which solutions are achieved. High-quality solutions minimize the number of control actions and avoid disruptive measures, such as load curtailment, whenever possible. Finally, runtime was recorded as the average wall-clock time, measured in seconds, required to reach a final solution.
Together, these metrics provide a comprehensive evaluation of the system’s performance, balancing effectiveness, efficiency, practicality, and solution robustness.
@@ V-D Experimental Results and Analysis
As shown in Fig.2a, the more capable and efficient models, gemini-2.5-flash and gpt-4.1-mini, achieved a perfect 100% success rate across all scenarios. The flagship models, gemini-2.5-pro and gpt-4.1, also performed robustly with an 100% success rate. The smaller models, gemini-2.5-flash-lite and gpt-4.1-nano, struggled with the more complex scenarios, achieving 63% and 44% success rates, respectively.
@@ VI-B Scalability Considerations
Furthermore, while this exploratory study has focused on the feasibility and strategic reasoning capabilities of Grid-Agent, a crucial direction for future work is the formal analysis of solution optimality. We plan to conduct rigorous benchmarking studies comparing the control actions proposed by Grid-Agent against traditional optimization techniques, particularly mixed-integer nonlinear Optimal Power Flow (OPF) solvers. Such a comparison will be essential to quantify the trade-offs between the LLM’s heuristic, rapid, and interpretable decision-making and the mathematically guaranteed optimality of conventional methods. This will help establish clear performance envelopes for where an agentic approach excels and where traditional optimizers remain indispensable, potentially leading to hybrid systems that leverage the strengths of both.
@@ VII Conclusion
Key achievements of this work include the development of a modular multi-agent architecture for decomposed problem-solving, an adaptive multi-scale representation framework that enables LLM application to larger networks, and a coordinated action optimization strategy that resolves multiple violations with minimal intervention. The integrated safety mechanisms, including sandboxed execution and automated rollbacks, ensure that all proposed solutions are validated before affecting the system, making the framework suitable for critical infrastructure applications.
@@ Acknowledgment
The authors acknowledge the support of the open-source community, particularly the developers of pandapower [18] and LangGraph frameworks that enabled this research.